OpenStudy (anonymous):

What values of x make the two expressions below equal? (2x+1)(x-7)/11(x-7)= 2x+1/11 A. All real numbers B. All real numbers except -1/2 C. All real numbers except 7 D. All real numbers except -1/2 and 7

OpenStudy (danjs):

missing parenthesis

OpenStudy (danjs):

maybe

OpenStudy (anonymous):

The best way to go about this problem is for you to plug in the numbers where stated "except"" to see if they work.

OpenStudy (danjs):

\[\frac{ (2x+1)(x-7) }{ 11(x-7) }=\frac{ 2x+1 }{ 11 }\] is that it

OpenStudy (anonymous):

yes that's what I wrote

OpenStudy (anonymous):

If not A, but if they do you'd be picking one of the B to D. Simple right?

OpenStudy (danjs):

group the entire numerator in a bra kets

OpenStudy (danjs):

simplify, (x-7) can cancel in the first, but you must keep in mind that x cant be 7, it will make the original expression undefined

OpenStudy (danjs):

then both are over 11, and you can multiply both sides by 11 to cancel those... left with 2x+1 = 2x+1 , true, all reals

OpenStudy (anonymous):

so if you cancel x-7 out, top and bottom you're left with (2x+1)-11 right?

OpenStudy (anonymous):

/11

OpenStudy (danjs):

yes, \[\frac{ 2x+1 }{ 11 }= \frac{ 2x+1 }{ 11 }\]

OpenStudy (danjs):

It is like saying, 1 = 1, True statement, any value for x is possible there

OpenStudy (danjs):

remember you canceled out (x-7) in the denominator though, so x-7 can not be zero

OpenStudy (anonymous):

same exact thing on both sides except the left one has parentheses

OpenStudy (danjs):

right, if you have the same things on both sides, then the variable can be any real number. Except here it can not be 7 since you cancelled out x-7 in the denominator, you cant have x-7=0 or x=7

OpenStudy (danjs):

zero in the bottom is bad

OpenStudy (anonymous):

so it would be all real numbers except 7?

OpenStudy (danjs):

yes

OpenStudy (anonymous):

thank you ;)-

OpenStudy (anonymous):

:)

OpenStudy (danjs):

i forget what they call those solutions x=7, some term

OpenStudy (anonymous):

did not know that myself haha

OpenStudy (danjs):

you can check, put x=7 and see if both sides are the same...

OpenStudy (anonymous):

i did

OpenStudy (danjs):

extraneous solutions, just popped into my head, check that definition out if it sounds familiar.